An accounting firm has three departments. Each department consists of account clerks, executives and accountants. The numbers are as follows;


Department 1

Accountants - 1

Executives - 2

Account clerks -.8


Department 2

Accountants - 3

Executives - 5

Account clerks - 14


Department 3 

Accountants - 2

Executives - 3

Account clerk - 12


If five workers are selected at random, find:


i) the number of ways at least three workers from Department 3 were chosen.


ii) the probability that two executives were chosen.


1
Expert's answer
2021-06-29T09:28:52-0400

i) The total number of workers in Department 1":1+2+8=11."

The total number of workers in Department 2":3+5+14=22."

The total number of workers in Department 1":2+3+12=17."


The number of workers in Department 1 and Department 2: "11+22=33"

The number of ways at least three workers from Department 3 were chosen is


"\dbinom{17}{3}\dbinom{33}{5-3}+\dbinom{17}{4}\dbinom{33}{5-4}+\dbinom{17}{5}\dbinom{33}{5-5}"

"=680(528)+2380(33)+6188(1)=443768"

(ii) The total number of workers: "11+22+17=50"

The total number of  executives: "2+5+3=10"

The probability that two executives were chosen is


"\dfrac{\dbinom{10}{2}\dbinom{50-10}{5-2}}{\dbinom{50}{5}}=\dfrac{45(9880)}{2118760}"

"=\dfrac{11115}{52969}\approx0.2098"